# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 7. Trustworthiness in Small Language Models
Holistic Evaluation of Language Models (HELM) (Liang et al., 2023) benchmarks a large number of LMs from various aspects, including a lot of metrics related to trustworthiness such as robustness and fairness. Do-Not-Answer (Wang et al., 2024d) introduces a dataset to evaluate how LMs act when they face content that should not be answered.  Wang et al. (2024d) also label the output of several LMs output on their dataset and then use the labeled data to train some classifiers. PromptRobust (Zhu et al., 2023c) constructs two kinds of adversarial prompts to evaluate LMs’ robustness: One kind is designed under non-adversarial settings with semantic integrity while another category is created under adversarial settings. Their results show that LMs perform poorly under such prompts. HaluEval (Li et al., 2023c) builds a dataset comprising both the samples generated by their proposed framework and human-labeled hallucinations. It facilitates analysis of when LMs produce hallucinated output and how well they detect hallucinated content. Then they use some strategies such as knowledge retrieval to help LMs better recognize hallucinations. Mo et al. (2024) evaluates the trustworthiness of open-source LMs, presenting a variety of scenarios such as fairness and privacy. Results show that smaller LMs sometimes outperform larger ones in terms of trustworthiness. PrivLM-Bench (Li et al., 2024b) is designed to evaluate the privacy issues in LMs. It enables a fair comparison of privacy-preserving LMs by considering more than just differential privacy parameters. FFT (Cui et al., 2023) introduces around two thousand carefully crafted examples to evaluate LMs’ performances on three trustworthy dimensions: factuality, fairness, and toxicity. Their results suggest that larger LMs do not always show better harmlessness. ROBBIE (Esiobu et al., 2023) first benchmarks various series of LMs using a lot of datasets, including two newly introduced datasets developed by ROBBIE. It also evaluates mitigation techniques designed to reduce bias and toxicity. TrustLLM (Sun et al., 2024b) is a comprehensive benchmark that contains a large number of datasets and various well-designed metrics to systematically evaluate various LMs across multiple trustworthy dimensions, including truthfulness, safety, fairness, robustness, privacy, and machine ethics. They also carefully design specific subcategories for each dimension. RAmBLA (Bolton et al., 2024a) evaluates the trustworthiness of four LMs as biomedical assistants from three dimensions: Robustness, High Recall, and Hallucination. RAmBLA suggests LMs with more parameters are less likely to cause hallucinations and may choose to reject providing an answer in uncertain situations. JailbreakBench (Chao et al., 2024) constructs a jailbreaking dataset named JBB-Behaviors and jailbreak artifacts to evaluate current LMs’ performance regarding jailbreaking. It also proposes a unified evaluation pipeline that can incorporate new jailbreak defense techniques. Xie et al. (2024b) tests online safety analysis methods, filling the gap where no methods focus on the generation phase. OR-Bench (Cui et al., 2024) constructs three datasets: OR-Bench-80K, OR-Bench-Hard-1K, and OR-Bench-Toxic, to systematically evaluate over-refusal problems in LMs, emphasizing the challenge of balancing safety alignment with the models’ usefulness. SORRY-Bench (Xie et al., 2024a) systematically tests 43 different LMs to see how they perform when facing requests that should be refused. They also collect more than annotations created by humans and find that fine-tuned 7B LMs can achieve performance comparable to GPT-4 scale LMs as evaluators. BeHonest (Chern et al., 2024) evaluates the honesty of LMs from three aspects: Self-Knowledge, Non-Deceptiveness, and Consistency. They use many different metrics for each aspect. For example, sycophancy rate and lying rate are adopted in Non-Deceptiveness. The results in both the Self-Knowledge and Consistency parts reveal that larger model sizes generally bring improved performance for the Llama-2 (Touvron et al., 2023b) and Llama-3 (Dubey et al., 2024) series. Hong et al. (2024) examines the effects of compression methods, including quantization and pruning, on the trustworthiness of language models. They find that pruning and extreme quantization significantly affect the trustworthiness of LMs. RUPBench (Wang and Zhao, 2024) comprises 15 reasoning datasets designed to assess the performance of LMs both in normal conditions and under various adversarial perturbations. Their results indicate that larger LMs generally demonstrate better resilience to perturbations.  Nakka et al. (2024) investigates the trust and ethical implications of SLMs deployed on personal devices. It reveals the vulnerabilities of on-device SLMs compared with their on-server counterparts.
@@
We summarize all the works above in Table 14. Please note that the dimensions listed in the table reflect only those relevant to our current focus; additional dimensions may be discussed in those works, but not included in the table. For example, TrustLLM (Sun et al., 2024b) also explores Machine Ethics.
@@ 8.1. Developing Efficient SLM Model Architecture
Although Transformers (Vaswani, 2017) are foundational in most language models, they face significant computational and memory challenges that worsen with model size, impacting training and autoregressive decoding. Recently, Mamba (Gu and Dao, 2023), a promising alternative, has emerged. It adapts the Multi-Head Attention mechanism to dynamically scale attention based on task demands, boosting efficiency, especially for simpler sequences. Mamba retains the content-aware capabilities of Transformers but scales linearly with input length, improving efficiency in both training and inference. Hence, developing domain-specific SLMs based on Mamba offers a promising research direction. Additionally, exploring other model architectures that enhance performance while reducing computational demands is also important.
@@ 8.4. Enhancing SLM Performance and Efficiency
  • Efficient Instruction Tuning of SLMs from LLMs-generated data. Enhancing the specialization of SLMs through instruction tuning from LLMs-generated data is crucial, yet finding the most cost-effective instructional strategies remains an underexplored status. Some key areas for exploration are: (1) Instruction Design Adaptability: The performance of LLMs and SLMs varies significantly with changes in instructions. Therefore, designing tailored instructions that effectively activate relevant sub-competencies and reasoning pathways in SLMs for specific tasks is crucial. This approach would optimize their ability to utilize instructional data, representing a significant future research direction. (2) SLM Capability Adaptability: Given that SLMs exhibit diverse capabilities across domains, simply supplying extensive data samples for instruction tuning is often inefficient, as SLMs may spend excessive time processing unnecessary data. To optimize efficiency when adapting to specific domains, we suggest first assessing the intrinsic capabilities of an SLM within those domains. Subsequently, one could select appropriate data and activate essential fine-grained capabilities to effectively adapt to domain shifts. This targeted approach ensures efficient and domain-specific instruction tuning. (3) Optimizing Data Efficiency: SLMs may possess missing or latent domain knowledge, and activating this latent knowledge may not require substantial data. Thus, identifying inherent knowledge within SLMs and determining the minimal data necessary for effective fine-tuning is a future direction. This research aims to optimize performance while minimizing training resources.
@@
8.5. Applications of SLMs: OOD, Personalization, Lifelong Learning
In real-world applications, SLMs often encounter out-of-distribution samples, need to provide personalized services, and need to be updated periodically to reflect new needs and new knowledge. Hence, there are several promising directions in terms of the real-world application of SLMs, which are listed as follows:
  • Enhancing OOD with Integrated RAG and Self-Adaptive Techniques. SLMs often encounter out-of-distribution (OOD) samples due to limited capabilities and training data. To enhance inference performance, Retrieval-Augmented Generation (RAG) and self-adaptive methods are utilized. RAG provides additional information, while self-adaptive methods generate pseudo labels for self-training. However, inaccuracies in pseudo-labels can reduce their effectiveness. A promising future direction is to integrate RAG with self-adaptive techniques to improve pseudo-label accuracy.
@@ 8.6. SLMs Assisting LLMs
In Section 6, we introduced existing works on the use of SLMs to assist LLMs. For instance, EFT (Mitchell et al., 2024) emulates fine-tuning on LLMs by leveraging behavior deltas between SLMs’ pre-trained and fine-tuned weights to alleviate the time-cost issues associated with fine-tuning LLMs; SlimPLM (Tan et al., 2024) detects missing knowledge in LLMs using a slim proxy SLM to accelerate knowledge injection; Contrastive Decoding (Li et al., 2023d) enhances text quality by maximizing the difference between the log probabilities of an expert LLM and an amateur SLM to mitigate issues of low-quality generation. The research on adopting SLMs to assist LLMs is still in its early stages, with many promising directions yet to be explored. We list some as follows:
  • Enhancing LLM Performance Across Broader Tasks Through SLM Integration. SLMs can outperform LLMs in certain scenarios. For example, SLMs often present fewer security vulnerabilities compared to their larger counterparts and demonstrate superior performance on easier samples in specific tasks (Li et al., 2024e; Ma et al., 2023a). Therefore, integrating SLMs with LLMs can promote the development of models that are not only more robust but also inherently safer. Currently, research in this domain is relatively sparse, suggesting that this collaborative framework could potentially be applied to a wider array of tasks.
@@
  • SLMs Assist in Managing Data Quality. LLMs tend to produce hallucinations and toxic content due to low-quality real-world training data. One solution is to remove these low-quality data (Wang et al., 2023b). However, directly eliminating low-quality content can diminish certain functionalities of LLMs, such as versatility (Wang et al., 2024b). Therefore, it is crucial to define more refined data quality assessment criteria across dimensions such as factuality, safety, and diversity (Wettig et al., 2024) for real-world data. Researching efficient data selection methods using small models represents a valuable research direction. Additionally, while synthetic data serves as a vital complement to scarce human-generated data (Long et al., 2024), the potential for small models to effectively manage synthetic data remains largely unexplored.
@@
  • SLMs Optimize Query and Reduce Noise for LLM RAG. For Retrieval-Augmented Generation (RAG) using LLMs, differing query requirements between LLMs and search engines pose a challenge. The query for LLMs is often abstract and difficult to handle by search engines, which require more detailed query keywords. Moreover, LLMs may not need all the information related to a query because they only require partial additional knowledge. Thus, intermediate agents are crucial to adapting LLM queries for search engines by clarifying the required detailed keywords that can search for necessary extra knowledge. Additionally, search engine outputs contain noises, requiring refinement to boost LLM efficiency. SLMs, skilled in a single task, are ideal for optimizing query rewriting and noise reduction in RAG systems, making their application in LLM RAG a promising research area.
  • Cloud LLM-Local SLM Collaboration. Cloud LLMs are commonly used, but privacy concerns limit them to processing all request aspects, necessitating the use of local SLMs to handle sensitive data (Zhang et al., 2024h). This setup allows for the specialized capabilities of SLMs to be leveraged on-device while still benefiting from the generalized functionalities of cloud-based LLMs. Researching optimal designs for collaboration that balance and enhance the capabilities of cloud-based LLMs and local SLMs is a crucial future direction.
  • Optimizing Specialized Training for Enhanced Model Performance. Specialized training and fine-tuning strategies are crucial for enhancing language model performance in specific tasks, such as coding. Models like DeepSeek-Coder illustrate the benefits of focused training, achieving significant gains in specialized domains. However, these models often struggle with broader or more complex requests, revealing a key trade-off. Future research could investigate integrating these specialized models with general SLMs and LLMs to merge specialization benefits with broader model versatility.
@@ 8.7. Trustworthy SLMs
  • Developing Trustworthy SLMs. In addition to understanding the trustworthiness of SLMs, developing trustworthy SLMs and improving their trustworthiness is also of critical importance. Essentially there are three directions worthy of investigation: (i) One promising direction is how to train trustworthy SLMs from scratch; (ii) As one can also obtain SLMs by compression LLMs, another promising direction is how to obtain trustworthy SLMs during compression or quantization of LLMs. In particular, if the LLM is trustworthy, how to preserve its trustworthiness during compression? If the LLM is non-trustworthy, how do we make SLM trustworthy during compression? (iii) Given a non-trustworthy SLM, how to fine-tune it to make it more robust?
@@ 9. Conclusion
This paper provides a comprehensive survey of SLMs with up to 7 billion parameters. We first clarify the definition of SLMs due to the current lack of clarity. We then outline foundational concepts essential for building SLMs. Subsequently, our review focuses on enhancement techniques such as knowledge distillation and quantization, along with strategies for adapting LLMs to SLM contexts. The subsequent section surveys representative SLMs, discussing their preferred datasets and architectural choices. We also examine their applications across various tasks as well as deployment strategies on devices and explore their role in augmenting the capabilities of LLMs. Another critical aspect discussed is their trustworthiness. The paper concludes with key insights intended to guide future research on small language models. Specifically, we highlight the following promising future directions.

### v2

@@ 8. Trustworthiness in Small Language Models
Holistic Evaluation of Language Models (HELM) (Liang et al., 2023) benchmarks a large number of LMs from various aspects, including a lot of metrics related to trustworthiness such as robustness and fairness. Do-Not-Answer (Wang et al., 2024f) introduces a dataset to evaluate how LMs act when they face content that should not be answered.  Wang et al. (2024f) also label the output of several LMs output on their dataset and then use the labeled data to train some classifiers. PromptRobust (Zhu et al., 2023c) constructs two kinds of adversarial prompts to evaluate LMs’ robustness: One kind is designed under non-adversarial settings with semantic integrity while another category is created under adversarial settings. Their results show that LMs perform poorly under such prompts. HaluEval (Li et al., 2023d) builds a dataset comprising both the samples generated by their proposed framework and human-labeled hallucinations. It facilitates analysis of when LMs produce hallucinated output and how well they detect hallucinated content. Then they use some strategies such as knowledge retrieval to help LMs better recognize hallucinations. Mo et al. (2024) evaluates the trustworthiness of open-source LMs, presenting a variety of scenarios such as fairness and privacy. Results show that smaller LMs sometimes outperform larger ones in terms of trustworthiness. PrivLM-Bench (Li et al., 2024c) is designed to evaluate the privacy issues in LMs. It enables a fair comparison of privacy-preserving LMs by considering more than just differential privacy parameters. FFT (Cui et al., 2023) introduces around two thousand crafted examples to evaluate LMs’ performances on three trustworthy dimensions: factuality, fairness, and toxicity. Their results suggest that larger LMs do not always show better harmlessness. ROBBIE (Esiobu et al., 2023) first benchmarks various series of LMs using a lot of datasets, including two newly introduced datasets developed by ROBBIE. It also evaluates mitigation techniques designed to reduce bias and toxicity. TrustLLM (Sun et al., 2024b) is a comprehensive benchmark that contains a large number of datasets and various well-designed metrics to systematically evaluate various LMs across multiple trustworthy dimensions, including truthfulness, safety, fairness, robustness, privacy, and machine ethics. They also carefully design specific subcategories for each dimension. RAmBLA (Bolton et al., 2024a) evaluates the trustworthiness of four LMs as biomedical assistants from three dimensions: Robustness, High Recall, and Hallucination. RAmBLA suggests LMs with more parameters are less likely to cause hallucinations and may choose to reject providing an answer in uncertain situations. JailbreakBench (Chao et al., 2024) constructs a jailbreaking dataset named JBB-Behaviors and jailbreak artifacts to evaluate current LMs’ performance regarding jailbreaking. It also proposes a unified evaluation pipeline that can incorporate new jailbreak defense techniques. Xie et al. (2024b) tests online safety analysis methods, filling the gap where no methods focus on the generation phase. OR-Bench (Cui et al., 2024) constructs three datasets: OR-Bench-80K, OR-Bench-Hard-1K, and OR-Bench-Toxic, to systematically evaluate over-refusal problems in LMs, emphasizing the challenge of balancing safety alignment with the models’ usefulness. SORRY-Bench (Xie et al., 2024a) systematically tests 43 different LMs to see how they perform when facing requests that should be refused. They also collect more than annotations created by humans and find that fine-tuned 7B LMs can achieve performance comparable to GPT-4 scale LMs as evaluators. BeHonest (Chern et al., 2024) evaluates the honesty of LMs from three aspects: Self-Knowledge, Non-Deceptiveness, and Consistency. They use many different metrics for each aspect. For example, sycophancy rate and lying rate are adopted in Non-Deceptiveness. The results in both the Self-Knowledge and Consistency parts reveal that larger model sizes generally bring improved performance for the Llama-2 (Touvron et al., 2023b) and Llama-3 (Dubey et al., 2024) series. Hong et al. (2024) examines the effects of compression methods, including quantization and pruning, on the trustworthiness of language models. They find that pruning and extreme quantization significantly affect the trustworthiness of LMs. RUPBench (Wang and Zhao, 2024) comprises 15 reasoning datasets designed to assess the performance of LMs both in normal conditions and under various adversarial perturbations. Their results indicate that larger LMs generally demonstrate better resilience to perturbations.  Nakka et al. (2024) investigates the trust and ethical implications of SLMs deployed on personal devices. It reveals the vulnerabilities of on-device SLMs compared with their on-server counterparts.
@@
Please note that the dimensions discussed in this section reflect only those relevant to our current focus; additional dimensions may be discussed in those works, but not listed in table 14. For example, TrustLLM (Sun et al., 2024b) also explores Machine Ethics.
@@ 9.1. Developing Efficient SLM Model Architecture
Although Transformers (Vaswani, 2017) are foundational in most language models, they face significant computational and memory challenges that worsen with model size, impacting training and autoregressive decoding. Recently, Mamba (Gu and Dao, 2023) has emerged as a promising alternative, adapting state space models to dynamically select inputs based on demands, thereby enhancing efficiency. Thereafter, xLSTM (Beck et al., 2024) demonstrates that an improved LSTM could function as an LLM, revealing the potential of traditional SSMs. The integration of global static information captured by SSMs with the dynamic information processing of Transformers could complement each other, leading to new architectures that balance effectiveness and efficiency.
9.2. Addressing SLM Training Inefficiencies
One study (Diehl Martinez et al., 2024) explores the disparate learning dynamics between small and large language models. Utilizing the Pythia model suite, the research demonstrates that layers’ activations in larger models converge more rapidly and monotonically to their final states. This phenomenon is associated with a higher proportional effective rank (PER) in the parameters and gradients of larger models. The analysis enhances our understanding of training inefficiencies in small models and provides insights for future efforts, such as developing methods to increase the PER of layers’ parameters.
@@ 9.5. Enhancing SLM Performance and Efficiency
  • Efficient Instruction Tuning of SLMs from LLMs-generated data. Enhancing the specialization of SLMs through instruction tuning from LLMs-generated data is crucial, yet finding the most cost-effective instructional strategies remains an underexplored status. Some key areas for exploration are:
    (1) Instruction Design Adaptability: The performance of LLMs and SLMs varies significantly with changes in instructions. Therefore, designing tailored instructions that effectively activate relevant sub-competencies and reasoning pathways in SLMs for specific tasks is crucial. This approach would optimize their ability to utilize instructional data, representing a significant future research direction.
    (2) SLM Capability Adaptability: Given that SLMs exhibit diverse capabilities across domains, simply supplying extensive data samples for instruction tuning is often inefficient, as SLMs may spend excessive time processing unnecessary data. To optimize efficiency when adapting to specific domains, we suggest first assessing the intrinsic capabilities of an SLM within those domains. Subsequently, one could select appropriate data and activate essential fine-grained capabilities to effectively adapt to domain shifts. This targeted approach ensures efficient and domain-specific instruction tuning.
    (3) Optimizing Data Efficiency: SLMs may possess missing or latent domain knowledge, and activating this latent knowledge may not require substantial data. Thus, identifying inherent knowledge within SLMs and determining the minimal data necessary for effective fine-tuning is a future direction. This research aims to optimize performance while minimizing training resources.
@@
9.6. Applications of SLMs
In real-world applications, SLMs often need to provide personalized services and need to be updated periodically to reflect new needs and new knowledge. Hence, there are several promising directions in terms of the real-world application of SLMs, which are listed as follows:
@@ 9.6. Applications of SLMs
  • Strategic Use of SLMs and LLMs in Multi-Agent Systems. LLMs can function as agents; however, their extensive capabilities are often underutilized in many scenarios, leading to resource wastage. Consequently, strategically routing to appropriately capable SLMs and LLMs within multi-agent systems can optimize cost and functionality.
9.7. Multimodal SLMs
Research on small language models also includes multimodal data. For example, SmolVLM (Face, 2024) is a compact model that handles image and text inputs to produce text outputs, suitable for on-device use and various multimodal tasks. SOLO (Chen et al., 2024b) integrates vision and language processing in a single 7B Transformer model. The limited scope of existing research on multimodal SLMs provides a compelling impetus for researchers to investigate the integration of various modalities, including audio and graphs.
@@ 9.8. SLMs Assisting LLMs
In Section 6, we introduced existing works on the use of SLMs to assist LLMs. For instance, EFT (Mitchell et al., 2024) emulates fine-tuning on LLMs by leveraging behavior deltas between SLMs’ pre-trained and fine-tuned weights to alleviate the time-cost issues associated with fine-tuning LLMs; SlimPLM (Tan et al., 2024) detects missing knowledge in LLMs using a slim proxy SLM to accelerate knowledge injection; Contrastive Decoding (Li et al., 2023e) enhances text quality by maximizing the difference between the log probabilities of an expert LLM and an amateur SLM to mitigate issues of low-quality generation. The research on adopting SLMs to assist LLMs is still in its early stages, with many promising directions yet to be explored. We list some as follows:
  • Enhancing LLM Performance Across Broader Tasks Through SLM Integration. SLMs can outperform LLMs in certain scenarios. For example, SLMs often present fewer security vulnerabilities compared to their larger counterparts and demonstrate superior performance on easier samples in specific tasks (Li et al., 2024f; Ma et al., 2023a). Therefore, integrating SLMs with LLMs can promote the development of models that are not only more robust but also inherently safer. Currently, research in this domain is relatively sparse, suggesting that this collaborative framework could potentially be applied to a wider array of tasks.
@@
  • SLMs Assist in Managing Data Quality. LLMs tend to produce hallucinations and toxic content due to low-quality real-world training data. One solution is to remove these low-quality data (Wang et al., 2023b). However, directly eliminating low-quality content can diminish certain functionalities of LLMs, such as versatility (Wang et al., 2024d). Therefore, it is crucial to define more refined data quality assessment criteria across dimensions such as factuality, safety, and diversity (Wettig et al., 2024) for real-world data. Researching efficient data selection methods using small models represents a valuable research direction. Additionally, while synthetic data serves as a vital complement to scarce human-generated data (Long et al., 2024), the potential for small models to effectively manage synthetic data remains largely unexplored.
@@
  • SLMs Optimize Query and Reduce Noise for LLM RAG. For Retrieval-Augmented Generation (RAG) using LLMs, differing query requirements between LLMs and search engines pose a challenge. The query for LLMs is often abstract and difficult for search engines to handle, so they require more detailed query keywords. Moreover, LLMs may not need all the information related to a query because they only require partial additional knowledge. Thus, intermediate agents are crucial to adapting LLM queries for search engines by clarifying the required detailed keywords that can search for necessary extra knowledge. Additionally, search engine outputs contain noises, requiring refinement to boost LLM efficiency. SLMs, skilled in a single task, are ideal for optimizing query rewriting and noise reduction in RAG systems, making their application in LLM RAG a promising research area.
  • SLMs safeguard LLM. Resources such as the Llama 2 Responsible Use Guide strongly advocate for the implementation of robust guardrails in products that utilize Generative AI. SLMs can be strategically designed to serve as such guardrails, mitigating risks associated with both inputs and outputs from the model. This approach ensures safeguards against the generation of high-risk or policy-violating content and provides protection against adversarial inputs and attempts to compromise the model. Future research can investigate the various safety roles that SLMs play in protecting LLMs.
9.9. Synergy between Small and Large Language Models
In Section 7, we discussed how small and large language models can complement each other. For example, CoGenesis (Zhang et al., 2024h) integrates SLMs for private data and LLMs for broader context, while Synergy of Thoughts (Shang et al., 2024) uses SLMs for initial reasoning and LLMs for conflict resolution. CROSSLM (Deng et al., 2023) shows how privacy can be preserved by training SLMs locally to support LLMs without data exposure. Research in this area is still evolving, and we outline several promising future directions below:
  • Refined Cloud-Edge Division of Labor. Current research mainly focuses on splitting tasks between edge-based SLMs and cloud-based LLMs along privacy-sensitive and non-sensitive data boundaries. A potential future direction involves more granular task partitioning: determining which subtasks should be handled locally by SLMs (e.g., initial data filtering, quick semantic parsing) and which should be delegated to the cloud-based LLM (e.g., advanced reasoning, complex generation). This approach can further optimize latency, privacy, and resource utilization.
  • Adaptive On-Device Specialization for Dynamic Environments. Although SLMs have shown the ability to handle private or personalized data locally, continuous changes in user preferences, application requirements, and data distributions pose challenges. Future work can explore adaptive strategies where edge-based SLMs dynamically specialize or update their parameters, guided by the cloud-based LLM. For instance, the LLM can periodically distill new knowledge into the SLM or provide feedback signals to help the SLM adapt to evolving scenarios.
@@ 9.10. Trustworthy SLMs
  • Developing Trustworthy SLMs. Developing trustworthy SLMs is crucial, with three key research directions: (i) Training SLMs to be trustworthy from scratch; (ii) Ensuring SLMs retain or gain trustworthiness when compressed from LLMs—maintaining trustworthiness if the LLM is trustworthy and instilling trustworthiness if it is not; (iii) Fine-tuning non-trustworthy SLMs to enhance their robustness.
@@ 10. Conclusion
This paper provides a comprehensive survey of Small Language Models (SLMs) with up to 7 billion parameters. Initially, we address the need to clearly define SLMs due to existing ambiguities in their characterization. We then present the foundational concepts essential for constructing SLMs. The survey progresses to explore enhancement techniques, including knowledge distillation and quantization, as well as strategies for adapting Large Language Models (LLMs) to SLM contexts. We survey representative SLMs, both general-domain and domain-specific, discussing their preferred datasets and architectural decisions. We also assess their applications across various tasks and deployment strategies on devices. Further, we investigate their role in augmenting the capabilities of LLMs, serving as proxies for fine-tuning and facilitating two types of synergies: cloud-local and task-centric. Additionally, we discuss the critical aspect of their trustworthiness. The paper concludes with key insights aimed at guiding future research on small language models.